OpenStudy (anonymous):

Find the area of a trapezoid whose parallel sides measure 8.5 m and 4.5 m, and whose altitude is 6.5 m.

OpenStudy (anonymous):

Area of a trapezoid = 1/2 h (sum of the two bases or the two parallel sides) Area = 1/2(6.5)(8.5 + 4.5) = 42.25 square meters. Note that the altitude is the height of the trapezoid.